Determine the amount of charge stored on either plate of a capacitor

(4x10-6 F) when connected across a 12 V battery.


1
Expert's answer
2022-05-24T13:56:35-0400
"q=CV\\
=4*10^{-6}*12=4.8*10^{-5}\:\rm C"